A rectangle’s perimeter and it’s area have the same numerical value. The width of the rectangle is 3 units. What it the length of the rectangle in units?

Found 2 solutions by jim_thompson5910, Boreal:
Answer by jim_thompson5910(35256)


L = length of rectangle = unknown
W = width of rectangle = 3 units
P = perimeter of rectangle
P = 2L+2W
P = 2L+2*3
P = 2L+6

A = area of rectangle
A = L*W
A = L*3
A = 3L

The area and perimeter are the same, so we can equate the expressions found to solve for L
A = P
3L = 2L+6
3L-2L = 6
L = 6

Answer: The length is 6 units

Check:
P = 2L+2W
P = 2*6+2*3
P = 12+6
P = 18
The perimeter is 18 units when we have a length of 6 units and width of 3 units.

A = L*W
A = 6*3
A = 18
The area is the same as the perimeter. The answer has been confirmed.

Answer by Boreal(15235)

area is LW=3*L
perimeter is 2L+2W=6+2L
those two are equal, so that 6+2L=3L
L=6 units ANSWER
Perimeter is 18 units 2*6 + 2*3
Area is 18 units 6*3